# Stratified distance space improves the efficiency of sequential samplers for approximate Bayesian computation

### v1

@@ 3 Stratified sampling ABC SMC
generate observations $y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}$. This is carried out as a part of the weighted ABC sampling at each iteration. The probability that predicted observations generated from
$p(y^{(t+1,t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}):=\int p(y\mid\theta^{(t+1,t)})p(\theta^{(t+1,t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})\mathsf{d}\theta^{(t+1,t)}$
can be estimated at iteration $t$ using the frequencies
@@
Figure 3: Example 1 continued. After the first ABC SMC iteration the particles are mutated and new set of synthetic observations are generated. We calculate the probabilities for each of the acceptance regions which are used to re-weight the ABC posterior sample weights.
@@
We illustrate the cumulative predictive probability weighting in Figure 3. At iteration $t$, our approximation to the posterior importance sampling distribution can be constructed from the posterior distribution $\{(\theta^{(t)}_{i},y^{(t)}_{i},w_{i}^{(t)})\}_{i=1}^{N}$ as
$\left\{\left(\theta^{(t)}_{i},y^{(t)}_{i},\hat{w}^{(t)}_{i},I_{i}^{(t)}\right)\right\}_{i=1}^{N}=\left\{\left(\theta_{i}^{(t)},y_{i}^{(t)},\frac{w_{i}^{(t)}\cdot\widehat{W}_{k}^{(t)}}{\sum_{j=1}^{N}w_{j}^{(t)}\cdot\widehat{W}_{k}^{(t)}},k\right):y_{i}^{(t)}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right\}_{i=1}^{N}.$ (21)
@@
10:    Find $k$ such that $y^{(t+1,t)}_{m}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$
@@ 4 Stopping rule
$P(y^{(t)}\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)})\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})$ $=C_{l,k}^{(t)},\quad l=1,\ldots,T.$ (22)
@@
as the quantified difference between the target and the current state. Note that by the definition we can start monitoring the convergence only after the second iteration as all the parameters on the first round are from the prior instead of the components of the partitioned approximate posterior.
@@
Figure 4: Example 1 continued. Upper figure bars illustrate the predicted probability mass functions (pmf) 22. At each iteration we calculate pmf defined by (22) and the corresponding KL divergence (23) that is reported on the bottom Figure.
@@ 5 Results
We assume that accuracy-wise all the methods will perform similarly as the main difference of the methods is the sampling strategy. We investigate the acceptance rates, and estimation accuracies in three examples which have been widely used in ABC literature for benchmarking methods. As estimation accuracy we report the medians of iteration-wise sample means in addition to the interval: median of the sample mean ±1.96$\pm$ 1.96$\cdot$ 1.96 ⋅ median of the sample standard deviation. In addition to accuracy we report the KL-divergence of the predicted acceptance region probabilities to investigate its usability as an early termination indicator. Acceptance rates and KL convergence monitoring quantity are reported as medians and interquartile ranges over the repetitions.
Figure 5: Acceptance rates by iteration round for the four compared methods in the three experiments with 2, 3 and 4 parameters as described in Sections 5.1 , 5.2 and 5.3. Acceptance rates are consistently higher for the proposed method in all experiments.
@@ 5.1 Banana shaped distribution
As divergence function we use Euclidean distance. Methods use the threshold sequence [,[$\infty$,[ ∞ , $100,$ $50,$ $20,$ $10,$ $5,$ $2,$ 1]1]1 ], and at each iteration the posteriors are approximated with 2000 samples. Acceptance rates provided by the proposed methods are significantly improved compared to local and global sampling strategies as seen in Figure 5. Even using only the proposed novel transition kernel (Stratified simple) the acceptance rate is better, but the stratified sampling strategy does improve the results markedly over almost all iterations. The estimation accuracy is slightly improved with the proposed methods and early termination convergence monitoring does indicate that the final iteration could be unnecessary. It is notable that the KL might be unreliable in the early iterations of the algorithm as indicated by the initial increase in the monitored divergence.
@@ 5.2 Lotka-Volterra
Acceptance rates are again significantly better with the proposed methods as reported in Figure 5. The estimation accuracy is better with the proposed methods, but the algorithms have not yet reached a stable state, but all methods should have improved performance when the threshold sequence would be continued.
@@
We use uniform prior distributions for the parameters
$k$ $\thicksim\mathsf{Unif}(0,2),$
@@
Euclidean distance is used as the divergence metric. All methods use a threshold sequence [,[$\infty$,[ ∞ , $100,$ $70,$ $50,$ $30,$ $27,$ $23,$ 20]20]20 ], and at each iteration the posteriors are approximated with 5000 samples. We observe similar results than on the previous experiments. Acceptance rates are significantly improved given the new proposed methods as illustrated in the Figure 5. From Figure 8 we see that the estimation accuracy of the proposed method is a little better compared to local and global. The early termination quantity does not indicate convergence within the iterations which can also be seen the posterior mean estimate evolution.
@@ 6 Conclusions
Through simulations we confirm the feasibility of the method and demonstrate that it achieves significantly improved acceptance rate than the currently most popular versions of ABC SMC algorithm. A restriction of the method however is the requirement of fixing the threshold sequence beforehand. Further work could thus be done to relax this requirement and toimprove the adaptability of the method.
@@ Acknowledgements
This research was supported by Research Council of Norway, grant no. 299941.

### v2

@@ 3 Stratified sampling ABC SMC
generate observations $y\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}$. This is carried out as a part of the weighted ABC sampling at each iteration. The distribution of the posterior predictive observations from the posterior at $t$th iteration that we will denote here as $y^{(t+1,t)}$ is
$p(y^{(t+1,t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}):=\int p(y^{(t+1,t)}\mid\theta^{(t+1,t)})p(\theta^{(t+1,t)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)})\mathsf{d}\theta^{(t+1,t)}$
can be estimated with the probability mass distribution of the discretised events $y^{\prime}\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}|y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. This is approximated using the frequencies
@@
Figure 3: Example 1 continued. After the first ABC SMC iteration the particles are mutated and new set of synthetic observations are generated. We calculate the probabilities for each posterior component of generating observations in each of the acceptance regions. These probabilities are cumulatively improved along with the iterations. In this case we observe that as expected, posterior component related to $\Omega_{0,1}$ has the largest probabilities of generating observations that will be accepted in any round. These component-wise predictive probabilities are then used to re-weight the ABC posterior sample weights.
@@
We illustrate the cumulative predictive probability weighting in Figure 3. Cumulative predictive probabilities improve along with the iterations as we obtain realisations of posterior predictive observations from different posterior components. We wish to re-balance the importance sampling distribution by giving more weight to particles that have been observed to generate samples within acceptance regions that will be included in the subsequent iterations of the algorithm. At iteration $t$, our approximation to the posterior importance sampling distribution can be constructed from the posterior distribution $\{(\theta^{(t)}_{i},y^{(t)}_{i},w_{i}^{(t)},I^{(t)}_{i}\}_{i=1}^{N}$, where $I^{(t)}_{i}\in\{1,\ldots,T\}$ indicates the corresponding partition of the observation space. For example, if $I^{(t)}_{i}=k$, then $y^{(t)}_{i}\in\Omega_{\left[\epsilon_{k+1},k\right)}$. The empirical importance sampling distribution is $\{(\theta^{(t)}_{i},y^{(t)}_{i},\hat{w}^{(t)}_{i},I^{(t)}_{i})\}_{i=1}^{N}$, where $\hat{w}^{(t)}_{i}$ is the re-balanced weight. It is calculated using the stratified observation space by first normalizing the weights within each strata, then weighting them with (25) and finally re-normalising the whole set, as
$\hat{w}^{(t)}_{i}\propto\frac{w_{i}^{(t)}}{\sum_{j:I^{(t)}_{j}=I^{(t)}_{i}}w_{j}^{(t)}}\cdot\widehat{W}_{I_{i}^{(t)}}^{(t)}.$ (26)
Based on the re-balanced weights we draw on average $\widehat{W}_{k}^{(t)}$ of the samples from $\{(\theta^{(t)}_{i},y^{(t)}_{i},w_{i}^{(t)},I^{(t)}_{i}):I_{i}^{(t)}=k\}_{i=1}^{N}$ because the predictive distribution conditioned on $y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$ have been approximated empirically to generate $\widehat{W}_{k}^{(t)}$ of the samples within $\Omega_{\left[0,\epsilon_{t+1}\right)}$. This stratified sampling approach should accelerate the acceptance rate especially in the early iterations of ABC SMC when the transition kernels can have large covariances. The largest weights for the samples calculated via (9) are assigned to the samples from the tails of the importance sampling distribution, when the corresponding synthetic observations are within the acceptance region. However, these outlying samples with large weights may also result in lower acceptance rate on the subsequent iterations.
We can estimate the locally optimal covariance kernel (22) based on the weighted sample by
$\hat{\Sigma}_{i}^{(k)}$ $=\sum_{j=1}^{N^{\prime}}\hat{\omega}^{(t)}_{I(j)}\left(\theta_{i}^{(t)}-\theta_{I(j)}^{(t)}\right)\left(\theta_{i}^{(t)}-\theta_{I(j)}^{(t)}\right)^{T},$
where $\theta^{(t)}_{i}$ is associated to acceptance region $\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$. The index set $\{I(1),\ldots,I(N^{\prime})\}\subset\{1,\ldots,N\}$ is such that $I^{(t)}_{I(i)}\geq k+1$, and the re-normalised weights are calculated as
$\hat{\omega}^{(t+1)}_{I(i)}=\frac{\hat{w}_{I(i)}^{(t)}}{\sum_{i=1}^{N^{\prime}}\hat{w}_{I(i)}^{(t)}}.$ (27)
@@
10:     Find $k$ such that $y^{(t+1,t)}\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}$
@@ 4 Stopping rule
$\mathsf{P}\left(y^{(t)}\in\Omega_{\left[\epsilon_{l+1},\epsilon_{l}\right)}\mid y\in\Omega_{\left[\epsilon_{k+1},\epsilon_{k}\right)}\right)$ $=C_{l,k}^{(t)},\quad l=1,\ldots,T.$ (28)
@@
as the quantified difference between the target and the current state. Note that by the definition we can start monitoring the convergence only after the second iteration as all the parameters on the first round are from the prior instead of the components of the partitioned approximate posterior. It is also notable that the early stopping rule monitors whether we have achieved posterior predictive performance of the final iteration round. Therefore if the final threshold is very small, it is possible that our empirical approximation of $C_{:,T}^{(t)}$ is poor, or in the worst case, does not contain a single sample. It is recommended to set up a minimum number of samples that are required for approximating $C_{:,T}^{(t)}$, and only after it is achieved, the early stopping rule monitoring is activated.
@@
Figure 4: Example 1 continued. Upper figure bars illustrate the predicted probability mass functions (pmf) (28). At each iteration we calculate pmf defined by (28) and the corresponding KL divergence (29) that is reported on the bottom figure.
@@ 5 Results
We assume that accuracy-wise all the methods will perform similarly as the main difference of the methods is the sampling strategy. We investigate the acceptance rates, and estimation accuracies in three examples which have been widely used in ABC literature for benchmarking methods. In addition, we investigate the performance of the proposed method when solving a real problem of inferring transmission dynamics of Streptococcus pneumoniae in day care centers. As estimation accuracy we report the medians of iteration-wise sample means in addition to the interval: median of the sample mean ±$\pm 1.96$cdot median of the sample standard deviation. In addition to accuracy we report the KL-divergence of the predicted acceptance region probabilities to investigate its usability as an early termination indicator. Acceptance rates, total number of generated samples and KL convergence monitoring quantity are reported as medians and interquartile ranges over the repetitions.
Figure 5: Acceptance rates by iteration round for the four compared methods in the three experiments with 2, 3 and 4 parameters as described in Sections 5.1 , 5.2 and 5.3. Acceptance rates are consistently higher for the proposed method in all experiments. The benefit of improved acceptance rate cumulates over the iteration rounds.
Figure 6: Total cumulative count of generated samples by iteration round for the four compared methods in the three experiments with 2, 3 and 4 parameters as described in Sections 5.1 , 5.2 and 5.3. The total cumulative number of samples indicate that the new stratified sampling method is more effective than the locally optimal ABC SMC. In banana shaped and G-and-k distribution examples, the locally optimal ABC SMC has reached sixth iteration, whereas the new proposed method is already on the final iteration round. On Lotka-Volterra model, when newly proposed method has reached the final iteration round, other methods have spent at least the same number of samples to reach the third iteration.
@@ 5.1 Banana shaped distribution
As divergence function we use Euclidean distance. Methods use the threshold sequence [,[$\infty$,[ ∞ , $100,$ $50,$ $20,$ $10,$ $5,$ $2,$ 1]1]1 ], and at each iteration the posteriors are approximated with 2000 samples. Acceptance rates provided by the proposed methods are significantly improved compared to local and global sampling strategies as seen in Figure 5. Even using only the proposed novel transition kernel (Stratified simple) the acceptance rate is better, but the stratified sampling strategy does improve the results markedly over almost all iterations. The estimation accuracy is slightly improved with the proposed methods and early termination convergence monitoring does indicate that the final iteration could be unnecessary. It is notable that the KL might be unreliable in the early iterations of the algorithm as indicated by the initial increase in the monitored divergence. The cumulative number of generated samples to achieve the required amount of samples is reported in Figure 6. Assuming the we would terminate the execution at the seventh round, we observe that e.g. the locally optimal ABC SMC has reached only the fourth iteration round given the same simulation budget. If we compare the estimation errors reported in Figure 7 on fourth and seventh iteration rounds of locally optimal ABC SMC and stratified sampling ABC SMC, respectively, we notice that the proposed method is much more efficient.
@@ 5.2 Lotka-Volterra
Acceptance rates are again significantly better with the proposed methods as reported in Figure 5. The estimation accuracy is better with the proposed methods, but the algorithms have not yet reached a stable state, but all methods should have improved performance when the threshold sequence would be continued as can be seen in Figure 8. The total number of generated samples is less than a half with the stratified sampling ABC SMC than with the other methods on the final iteration round as reported on Figure 6.
@@
Euclidean distance is used as the divergence metric. All methods use a threshold sequence [,[$\infty$,[ ∞ , $100,$ $70,$ $50,$ $30,$ $27,$ $23,$ 20]20]20 ], and at each iteration the posteriors are approximated with 5000 samples. We observe similar results than on the previous experiments. Acceptance rates are significantly improved given the new proposed methods as illustrated in the Figure 5. From Figure 9 we see that the estimation accuracy of the proposed method is a little better compared to local and global. The early termination quantity does not indicate convergence within the iterations which can also be seen the posterior mean estimate evolution. In Figure 6 we see that the stratified method is a couple of iterations ahead simulation-budget-wise to locally optimal ABC SMC.
@@
6 Bacterial transmissions in day care centers
We investigate a problem of estimating the transmission dynamics of Streptococcus pneumoniae from strain prevalence data collected in day care units in Oslo metropolitan area. Originally, the transmission model parameters were solved using ABC SMC in the study by Numminen et al. (2013). Our formulation of the problem follows a slightly modified model and the set of summary statistics following (Gutmann and Corander, 2016). We use a model where the number of the observed children equals the size of the center, i.e. we don’t simulate unobserved children. The data described in detail in Vestrheim et al. (2008) consists of samples collected from 611 healthy children from 29 different day care centers. The 33 most prevalent strains were included in the data set.
The transmission dynamics are modeled separately for each day care center as a continuous time Markov process
$\mathsf{P}\left(I_{is}(t+\delta t)=1\mid I_{is}(t)=0\right)=\beta E_{s}(I(t))+\Lambda P^{s}+o(\delta t),\text{ if }\sum_{j=1}^{N^{s}}I_{ij}(t)=0,$ (37)
$\mathsf{P}\left(I_{is}(t+\delta t)=1\mid I_{is}(t)=0\right)=\theta(\beta E_{s}(I(t))+\Lambda P^{s}+o(\delta t)),\text{ if }\sum_{j=1}^{N^{s}}I_{ij}(t)>0,I_{is}=0$ (38)
$\mathsf{P}\left(I_{is}(t+\delta t)=0\mid I_{is}(t)=1\right)=\gamma+o(\delta t),$ (39)
where $\beta\thicksim\mathsf{Unif}(0,11),\Lambda\thicksim\mathsf{Unif}(0,2)$ and $\theta\thicksim\mathsf{Unif}(0,1)$ are the internal infection parameter, the external infection parameter and the co-infection parameter, respectively. The first equation defines the probability for a child $i$ to be colonised with strain $s$ when the individual is not colonised with any of the other strains in the model, while the second equation described the probability when the child is already being colonised with at least one other strain. The last equation describes the recovery rate from colonization. We follow the formulation in Numminen et al. (2013) and set $\gamma=1$, and all other variables are scaled accordingly. The data is a single time point snapshot of the colonizations. We set the simulation time to $T=10$ time units, and use the state in the end as our observation. The resulting data consists of $N^{\text{dcc}}$ binary matrices $I_{is}^{(n)},i=1,\ldots,N^{c,\text{dcc}},s=1,\ldots,N^{s},n=1,\ldots,N^{\text{dcc}}$, where $N^{c,\text{dcc}}$ is the number of children in the day care center $n$, $N^{s}$ is the number of bacterial strains and $N^{\text{dcc}}$ is the number of day care centers.
We use a set of four summary statistics for each of the day care centers, i.e. Shannon index of of the strains, the number of different strains, the prevalence of carriage among children and the prevalence of more than one colonization among the children. The summarised data of size 4 from 29 day care centers was first scaled based on the observed data so that the maximum of each of the summaries would be one. After the summary statistic scaling, the empirical cdfs of the joint set of 116 summaries was then compared via $L_{1}$-norm scaled with the coefficient 1/116.
The two versions of the stratified sampling approach are tested along with the locally optimal ABC SMC to infer the model parameters with $5000$ approximate posterior samples using a threshold sequence $[\infty,0.5,0.1,0.08,0.05]$. The results illustrated in Figure 10 indicate that there are no significant differences in the inferred approximate posterior. Using the proposed algorithm, we do obtain the final posteriors by using less than half of the simulations when comparing the stratified and the locally optimal versions in Figure 11. There is a discrepancy between the posterior inferred by Numminen et al. (2013) and our approximate posteriors in the estimate of the parameter $\Lambda$ that is most likely caused by the difference in the used models, as described earlier in the section.
Figure 10: Evolution of the sample mean and the $95\%$-credible interval of the three parameters of the day care center model described in Section 6. The horizontal dashed lines are the sample mean and $95\%$-credible intervals as reported in Numminen et al. (2013).
Figure 11: Acceptance rates and the total number of generated samples by iteration round for the three compared methods in the day care center model described in Section 6. Acceptance rates are consistently higher for the proposed method in all experiments. The benefit of improved acceptance rate cumulates over the iteration rounds, and eventually the new proposed method requires only half of the simulator queries to reach the final approximate posterior.
@@ 7 Conclusions
Our simulations confirm feasibility of the proposed method and demonstrate that it achieves significantly improved acceptance rate than the currently most popular versions of the ABC SMC algorithm. Even if there are no large differences between the inference methods in estimation accuracy when considering specific iteration rounds, the improved acceptance rate means that the newly proposed method achieves the accuracy levels earlier. A restriction of the method, however, is the requirement of determining the threshold sequence beforehand. In practice this means that it is necessary to obtain problem domain specific knowledge about the behavior of the discrepancy measure using explorative analyses of the simulator output with feasible inputs. Of note, many such pre-inference analyses would also match the good practices from the traditional Bayesian workflow (Gelman et al., 2020). In situations where the a priori analyses would not be considered feasible, it would be necessary to resort to adaptive algorithms not necessitating pre-specified thresholds. Many adaptive ABC SMC algorithms have been designed to effectively skip unnecessary evaluations of ABC posteriors by finding the threshold values based on the current posterior sample. Even if these methods are not designed to improve the acceptance rate, they can achieve good performance with less total evaluations by skipping unnecessary steps when using fixed threshold sequencing. Further work could thus be done to combine more efficient sampling with adaptation of the acceptance regions to the data.
@@ Acknowledgements
This research was supported by Research Council of Norway, grant no. 299941 and through its Centre of Excellence Integreat - The Norwegian Centre for knowledge-driven machine learning, project number 332645.